Determine the equilibrium concentration of the base NH3 given that pKb=4.76 and the equilibrium concentrations of OH− and NH+4 are 6.5×10−5 M, and 0.42 M respectively.

  • Report your answer with two significant figures.
Answer #1

ENHolea pkb – 4.76 (given) = -log kb = 4.76 u log kb = -4.76 -476 .: Ky= 10 Now, reaction is: NHg69) + H20 hl) = NHT @q) + of